Give first the step you will use to separate the variable and then solve the equation:
a/5 = 7/15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Now we have to multiply both sides of the equation by 5,
Then we get,
= (a/5) × 5 = (7/15) × 5
= a = 7/3

Thanks